# Quaffure: Real-Time Quasi-Static Neural Hair Simulation

### v1

@@ 1 Introduction
In this paper, we propose a novel and efficient neural hair simulation approach to model hair quasi-statics. Our method leverages a learning-based method to predict physically plausible hair deformations, while also generalizing to various body poses, shapes, and hairstyles. We opted for training in a self-supervised manner, eliminating the need for expensive data generation and storage. Our approach achieves an inference performance of just a few milliseconds on standard consumer hardware, without requiring elaborate code optimizations or specialized hardware features. In contrast to physics-based simulation methods, the inference time of our approach is fixed and is independent of the number of strands in the groom or number of collisions in the scene. Our results are stable and the method scales to predicting hundreds of grooms simultaneously at interactive rates. In summary, our contributions are
@@
  • We demonstrate how a singled trained network efficiently generalizes to different hairstyles, body poses and shapes.
@@ 3.5 Physics-Based Self-Supervision
Although this full Cosserat rod model, which optimizes for both position and orientation offsets, can enable realistic modeling of hair, we observed impractically slow training times and in combination with slow convergence, this approach was deemed infeasible. Instead, we propose an alternative formulation, which optimizes for position offsets only with the following modified stretch-shear potential:
[…]
where $\mathbf{\tilde{\Gamma}}=\frac{\textbf{x}_{i+1}-\textbf{x}_{i}}{l_{\text{rest}}}-\textbf{d}_{3}$, with $\textbf{d}_{3}$ being the unit director along the edge, computed from the rigidly transformed groom. As the modified stretch-shear potential penalizes stretching and shearing simultaneously, we observed that hair would become overly rigid when a large stiffness value $\texttt{k}_{\text{Cosserat}}$ is used (which is common as hair is almost inextensible). Therefore, to relax $\texttt{k}_{\text{Cosserat}}$ for more natural hair deformation, we model resistance to stretching explicitly by adding a Hookean potential to penalize the length $l$ of each hair segment to be close to $l_{\text{rest}}$:
@@
Our final elastic potential include only two terms and we observe orders of magnitude faster training time when compared with the full Cosserat rod model
[…]
@@ 4 Experiments
Implementation Details: In all of our results, strands consist of $N=24$ vertices, which are encoded into $64\times 64$ texture maps. This choice of texture dimensions provides us with the ability to model several thousands of hair strands, which is sufficient for guide hair simulation. Prior work [13] shows high quality upsampling results for as little as 128 guide strands. We implement our method with PyTorch [28], with training and inference measurements performed on an AMD Ryzen Threadripper PRO 3975WX CPU and a single NVIDIA RTX A6000 GPU. We refer the reader to the supplementary material for additional details.
@@ 4.1 Qualitative Comparisons
Body Pose and Shape Generalization: Our deformation decoder network is able to ingest body shape parameters directly and effectively learns to adapt the deformations of the hair strands to produce intersection free results for a large variety of body shapes. Figure 4 showcases a large sample of pose variations for a diverse set of grooms while in Figure 5 we demonstrate body shape generalization.
@@
Table 1: Quantitative Comparisons. We compare our method to several optimization and simulation baselines as well as related work. Our method displays comparable metrics to directly optimizing for the positions but does so orders of magnitude faster. Compared to the state-of-the-art method GroomGen, our method produces better quality metrics while being generalizable to grooms, body shape and pose, with a comparable compute performance. We demonstrate that our approach provides the best trade-off between speed and quality.
@@
in seconds Intersection Preservation Preservation
Adam 179.38 0.22 103.53 76.15
L-BFGS 281.18 0.22 89.53 70.22
XPBD (GPU) [25] 63.26 0.01 57.96 18.10
@@ 4.2 Quantitative Comparisons
Comparisons: The quantitative results are listed in Table 1. As expected, directly modifying the simulated hair positions, either using Adam, L-BFGS or XPBD simulation to minimize the loss produces the highest quality results but at a high computational cost. Neural methods like GroomGen and ours are orders of magnitude faster. Our method produces significantly better quantitative results when compared to GroomGen at preserving strand lengths as well as the the preservation of the hair shape measured from the segment orientations. Furthermore, it is important to note that, unlike GroomGen, our method handles arbitrary motion and generalizes to body shape. Figure 8 shows that GroomGen results display frequent intersections of the hair strands with the body. These results demonstrate that our method provides the best trade-offs in terms of performance and quantitative results with the significant added benefit that we do not require any simulated training data.
@@
XPBD - Physics-based Simulation: We compare to the high-performance simulation method XPBD [25] since it is highly effective at exploiting GPU parallelism, which results in fast computation times. Physics-based simulation has been the de facto solution to hair animation for years due to the high realism, albeit at a high computational cost. Additionally, time integration of the physical system can often lead to instabilities, which will result in simulation failures or undesirable results. In contrast, our method does not display any instabilities. In our comparison, we model hair using the same energy as described in Subsection 3.5. To obtain quasi-static simulation results, we reset the velocities for each time step in the XPBD simulation loop, before applying gravity. We then integrate the positions and use $10000$ iterations to minimize the loss. We then run 30 time steps to obtain result close to convergence.
@@ 4.3 Performance
Predicting a single groom takes only $2.86$ ms and predicting a thousand at once requires just $0.3$ seconds. The performance is independent of the number of strands due to the fixed size of the decoder output. We plot performance for a varying number of groom and pose variations in Figure 9 and find that our algorithm scales linearly with the number of grooms. Strand count for the varying grooms used in training varies between 1500 and 2500.
@@ 4.4 Ablations & Discussion
We perform an ablation experiment to demonstrate the importance of our proposed Cosserat energy. Figure 10 showcases that our model is highly effective at maintaining hairstyle and curliness. We compare our results to those obtained by implementing a mass-spring energy [33] in our proposed framework. Our Cosserat model captures and maintains the intended hairstyle well without any manual intervention whereas the mass-spring model requires extensive parameter tuning to obtain satisfactory results and the hair still falls mostly flat under gravity. Our Cosserat model provides us with artistic control over the final result by manipulating the stiffness parameters during training. Figure 11 demonstrates how a lower Cosserat stiffness will result in a hairstyle that doesn’t preserve curliness as well as the variant with higher stiffness.
Limitations and Future Work: Our pose-based groom transformation module often produces hair where strands are partially initialized inside the body geometry. A better transformation model which results in fewer self-intersections would make it easier for the model to learn the required deformations to reach a quasi-static state. Nonetheless, we observe that our network is able to robustly deal with such cases. Current results are obtained using a single set of physical parameters describing the hair material properties. This is a major advantage as it shows that parameter tuning is not required to produce visually pleasing results. However, we plan to condition the network on material parameters to enable artistic control at inference.

### v2

@@ 1 Introduction
In this paper, we propose a novel and efficient neural hair simulation approach to model hair quasi-statics. Our method leverages a learning-based method to predict physically plausible hair deformations, while also generalizing to various body poses and shapes. We opted for training in a self-supervised manner, eliminating the need for expensive data generation and storage. Our approach achieves an inference performance of just a few milliseconds on standard consumer hardware, without requiring elaborate code optimizations or specialized hardware features. In contrast to physics-based simulation methods, the inference time of our approach is fixed and is independent of the number of strands in the groom or number of collisions in the scene. Our results are stable and the method scales to predicting hundreds of grooms simultaneously at interactive rates. In summary, our contributions are
@@
  • We demonstrate how a single trained network efficiently models different hairstyles, generalizing to body poses and shapes.
@@ 3.5 Physics-Based Self-Supervision
Although this full Cosserat rod model, which optimizes for both position and orientation offsets, can enable realistic modeling of hair, we observed impractically slow training times and in combination with slow convergence, this approach was deemed infeasible. Instead, we propose an alternative formulation, which optimizes for position offsets only with the following modified Cosserat potential:
$\mathcal{L}_{\text{Cosserat}}=\frac{1}{2}\sum_{\text{edges}}\texttt{k}_{\text{Cosserat}}\mathbf{\tilde{\Gamma}}^{T}\mathbf{\tilde{\Gamma}},$ (6)
where $\mathbf{\tilde{\Gamma}}=\frac{\textbf{x}_{i+1}-\textbf{x}_{i}}{l_{\text{rest}}}-\textbf{d}_{3}$, with $\textbf{d}_{3}$ being the unit director along the edge, computed from the rigidly transformed groom. As the modified Cosserat potential penalizes stretching and shearing simultaneously, we observed that hair would become overly rigid when a large stiffness value $\texttt{k}_{\text{Cosserat}}$ is used (which is common as hair is almost inextensible). Therefore, to relax $\texttt{k}_{\text{Cosserat}}$ for more natural hair deformation, we model resistance to stretching explicitly by adding a Hookean potential to penalize the length $l$ of each hair segment to be close to $l_{\text{rest}}$:
@@
Alternatively, one can choose to use the potential that penalizes shearing only: $\mathbf{\tilde{\Gamma}}=\frac{\textbf{x}_{i+1}-\textbf{x}_{i}}{l}-\textbf{d}_{3}$, with $l$ being the segment length. In practice, we observed similar results from both potentials. Our final elastic potential include only two terms and we observe orders of magnitude faster training time when compared with the full Cosserat rod model
$\mathcal{L}_{\text{elastic\_potential}}=\mathcal{L}_{\text{stretch}}+\mathcal{L}_{\text{Cosserat}}.$ (8)
@@ 4 Experiments
Implementation Details: In all of our results, strands consist of $N=24$ vertices, which are encoded into $64\times 64$ texture maps. This choice of texture dimensions provides us with the ability to model several thousands of hair strands, which is sufficient for guide hair simulation. We implement our method with PyTorch [28], with training and inference measurements performed on an AMD Ryzen Threadripper PRO 3975WX CPU and a single NVIDIA RTX A6000 GPU. We refer the reader to the supplementary material for additional details.
@@ 4.1 Qualitative Comparisons
Body Pose and Shape Generalization: Our deformation decoder network is able to ingest body shape parameters directly and effectively learns to adapt the deformations of the hair strands to produce intersection free results for a large variety of body shapes. Figure 4 showcases a large sample of pose variations for a diverse set of grooms while in Figure 5 we demonstrate body shape generalization. Implausible deformations are possible when deforming the body in non-physical ways, e.g. when bending the neck backwards beyond limits feasible in real-life and thus absent from our training data set.
@@
Table 1: Quantitative Comparisons. We compare our method to several optimization and simulation baselines as well as related work. Our method displays comparable metrics to directly optimizing for the positions but does so orders of magnitude faster. Compared to the state-of-the-art method GroomGen, our method produces better quality metrics while being generalizable to body shape and pose, with comparable compute performance, providing the best trade-off between speed and quality. Gravity potentials are provided for reference.
@@
in seconds Intersection Preservation Preservation Potential
Adam 179.38 0.22 103.53 76.15 3011.33
L-BFGS 281.18 0.22 89.53 70.22 3047.53
XPBD (GPU) [25] 63.26 0.01 57.96 18.10 3507.53
@@ 4.2 Quantitative Comparisons
Comparisons: The quantitative results are listed in Table 1. As expected, directly modifying the simulated hair positions, either using Adam, L-BFGS or XPBD simulation to minimize the loss produces the highest quality results but at a high computational cost. Neural methods like GroomGen and ours are orders of magnitude faster. Our method produces significantly better quantitative results when compared to GroomGen at preserving strand lengths as well as the preservation of the hair shape measured from the segment orientations. Furthermore, it is important to note that, unlike GroomGen, our method handles arbitrary motion and generalizes to body shape. Figure 8 shows that GroomGen results display frequent intersections of the hair strands with the body. These results demonstrate that our method provides the best trade-offs in terms of performance and quantitative results with the significant added benefit that we do not require any simulated training data.
@@
XPBD - Physics-based Simulation: The XPBD method [25] effectively exploits GPU parallelism, which results in fast computation times. Simulation has been the de facto solution to hair animation for years due to the high realism, albeit at a high computational cost. Additionally, time integration of the system can often lead to instabilities, which will result in failures or undesirable results. In contrast, our method does not display any time integration instabilities. In our comparison, we model hair using the same energy as described in Subsection 3.5. To obtain quasi-static simulation results, we reset the velocities for each time step in the XPBD simulation loop, before applying gravity. We integrate the positions and use $10000$ iterations to minimize the loss. We run 30 time steps to obtain results close to convergence.
@@ 4.3 Performance
Predicting a single groom takes $2.86$ ms and predicting a thousand at once requires $0.3$ seconds. The performance is independent of the number of strands due to the fixed size of the representation. Performance for a varying number of groom and pose variations is shown in Figure 9. Strand count for the varying grooms used in training varies between 1500 and 2500.
@@ 4.4 Ablations & Discussion
We perform an ablation experiment on the importance of our proposed Cosserat energy. Figure 10 shows that our model is highly effective at maintaining hairstyle and curliness. We compare our results to those using a mass-spring energy [33]. Our model maintains the intended hairstyle well without manual intervention whereas the mass-spring model requires extensive parameter tuning to obtain satisfactory results and the hair still falls mostly flat under gravity. Our model provides artistic control by manipulating the stiffness parameters. Figure 11 demonstrates the effect of varying material stiffness.
Limitations and Future Work: Our pose-based groom transformation module can produce hair where strands are partially initialized inside the body. A better model which results in fewer intersections would make it easier for the model to learn the required deformations. Nonetheless, we observe that our network is able to robustly handle this. Current results use a single set of physical parameters for the hair material properties. This shows that parameter tuning is not required to produce visually pleasing results. However, we plan to condition the network on material parameters to enable artistic control at inference.